OpenStudy (amorfide):

log[2](27) is the first term of a geometric sequence common ration is log[2](y) what are the ranges of the values of y to make a possible sum of series to infinity? what is the exact value of y if the infinity sum of series is 3? by sum of series to infinity i mean a/1-r just incase i worded it wrong and you didn't know what i meant.

OpenStudy (experimentx):

y > 2

OpenStudy (experimentx):

Solve for y \[ \log_2 27 \left ( \frac{1}{1 - \log_2y}\right ) = 3 \]

OpenStudy (experimentx):

better done this way \[ 1 = \log_2 2\] \[ 1 - \log_2y = \log_2(2/y)\] \[ \frac{\log_227}{\log_2(2/y)} = \log_{2/y} 27\] Now it seems lot easy to solve!!

OpenStudy (amorfide):

do not understand it seems you missed a step or two